Perform the indicated operations.
(4x-7y)(3x+2y)

Answer by rapaljer(4671)

(4x-7y)(3x+2y)

This is a FOIL problem. Do you know what that means? If not, send me an Email and I'll send you a couple of links.

F= First*First = 4x*3x=12x^2
O= Outer*Outer = 4x*2y=8xy
I=Inner *Inner = -7y*3x=-21xy
L=Last*Last = -7y*2y =-14y^2

Add them up and combine like terms:
12x^2 +8xy-21xy -14y^2
12x^2 -13xy - 14y^2

Dr. Robert J. Rapalje, Retired